# Contextual Prompt Learning for Vision-Language Understanding

### v1

@@ 4 Experimental Evaluation
We follow Zhou et al. [63] to evaluate our model on $11$ image classification dataset of varying complexity. The datasets include: generic classification datasets like ImageNet [6] and Caltech-101 [12]; curated fine-grained datasets like OxfordPets [47], StanfordCars [30], Flowers102 [46], Food101 [2] and FGVCAircraft [45]; scene, action, texture and satellite image recognition datasets from SUN397 [61], UCF101 [56], DTD [4] and EuroSat [23] respectively.
For the few-shot experiments, we follow Zhou et al. [62] to randomly sample datapoints for training. The models are evaluated on the entire test set to report the accuracy. For setting $(i)$ and $(ii)$ (inter-dataset and intra-dataset experiments), we use $16$-shots, consistent with the baseline methods [62]. We use the standard classification accuracy as the evalution metric across all experiments.
@@
We build our method on top of the publicly available implementation of CoCoOp [63]. For the CLIP architecture, we use ViT-B/$16$ as image encoder. For all our experiments, we use prompt token length of $4$. All our models are trained with a batch size $1$ for $10$ epochs on a single GPU system. Our models are trained on a $15$ GB Tesla T4444 GPU. We used the standard stochastic gradient descent (SGD) [52] as our optimizer. Our starting learning rate is $0.002$ and we use cosine learning rate scheduler. Our warm-up with a constant learning rate is $0.00001$.
@@
We compare our approach with the current state-of-the-art method CoCoOp [62], CoOp [63]and also large-scale zero-shot methodology CLIP [51]. While comparing with CLIP, we indeed compare our learned prompt embeddings with manually designed prompts. We closely follow most of our experimental setting with that of Zhou et al., [62].
@@
Table 2: Zero-shot image recognition results of CoPL across dataset. We trained models on caltech101 dataset and tested on other datasets.
4.1 Knowledge Transfer to Unseen Classes
One of the main issues of CoOp is that it is unable to generalize to the unseen classes while being trained on the base classes. In CoCoOp, the authors claim to bridge the gap by learning prompt conditioned on the global features of the image. But they miss out on capturing the local features of the image and do not give weightage to the semantically aligned learnable prompts, which is the key focus area of our work. Following the implementation of Zhou et al., [62], we conducted our experiments on the above mentioned $11$ datasets both for seen and unseen classes. While training is only conducted on the base classes, during testing we transfer the learnt knowledge to classify unseen classes as well as seen classes.
From Table 1, we can see that across datasets, there is a decrease in performance for CoOp methodology over the zero-shot large-scale method CLIP. Though there is an improvement for the seen classes, the failure across unseen classes makes the model less effective compared to CLIP. Though introduction of conditional prompting improved the performance of CoCoOp, in most of the cases it fails to generalize to unseen classes with the dynamically learnt prompts (in $7$ out of $11$ datasets CLIP has outperformed CoCoOp methodology). This raises the question of how to effectively learn the prompts which are more generalizable and align with the input image features?
Interestingly, our proposed approach CoPL, improves the accuracy of unseen classes over CoCoOp and CoOp for $10$ different tasks. Moreover, the harmonic mean for all these $10$ tasks are greater than the current state-of-the-art model, which specifies that the learned prompts are more generalizable across domains and tasks. It is important to note that CoPL comfortably outperforms manual prompt based methodology CLIP on $5$ different tasks. With the relatively complex dataset like EuroSAT, where localization is hard to be aligned due to the nature of satellite imagery, for the seen classes CoPL outperformed CoCoOp by $1.8\%$ and CLIP by $32.8\%$ on accuracy.
@@ 4.1 Knowledge Transfer to Unseen Classes
In Figure 3, we further analyse the absolute performance gain on the unseen classes over CoOp and CoCoOp methodologies. Interestingly, when compared with CoOp, for the dataset UFC101, CoPL improve the accuracy by $20.6\%$, enhancing the capability of the dynamic prompt vectors to learn semantic relatedness conditioned on the local image features. For the FGVCAircraft dataset, the absolute performance gain of $9.0\%$ and $7.6\%$ for CoOp and CoCoOp respectively highlights the efficacy of the model. Performance boost of $8.9\%$ on the DTD dataset, confirms the capability of COPL in identifying different textures, opening up the possibility to deploy the model for design understanding. The average performance gain of $2.7\%$ on the FGVCAircraft dataset over CLIP, defines the ability of CoPL to transfer the learnt knowledge to the relatively complex unseen classes.
@@
In general, while considering both seen and unseen dataset, CoPL outperforms CoCoOp by $2.7\%$ in accuracy. At the same time, it achieves a gain of $6.8\%$ on accuracy over manual prompt base methodology CLIP, suggesting that conditional alignment of prompts to local features of the image can generalize the model better to diverse recognition tasks. Ideally, the manually designed prompts should have been well aligned with the images, thus CLIP can be considered as the best performing model. It is interesting to note that CoPL, on an average across all $11$ datasets, achieves state-of-the-art performance by outperforming CLIP by $1.4\%$ in accuracy. This justifies our argument of making dynamic prompt vectors more contextualized like human annotated prompts by observing the different local features of the image.
4.2 Inter Dataset Zero-shot Performance
Here, we test the mettle of our approach on an even more challenging setting. We learn the model on a dataset, and evaluate the model on a different dataset to see how transferable the learned representations are.
@@ 4.2 Inter Dataset Zero-shot Performance
Concretely, we train the model on Caltech101 dataset and evalute its performance on the rest of the datasets. Observe, in Table 2, CoPL outperforms CoCoOp across most of the datasets. For complex dataset like FGVCAircraft, CoPL gains $2.9\%$ more accuracy without being explicitly trained on this dataset. As the Caltech101 is a general purpose object classification dataset, we observe impressive performance on OxfordPets and Food101 dataset. Interestingly, though trained on object classification dataset, CoPL is able to transfer the knowledge to do texture recognition task on DTD dataset, outperforming CoCoOp by $2.5\%$ on accuracy. In the most challenging EuroSAT dataset, which has very different image sets (consists of images taken from satellite), CoPL achieves $58.1\%$ accuracy which is comparable to the performance of the CoCoOp. These highlights that CoPL able to learn the semantics of the images by identifying the localized features and align it with the learnable textual prompts.
4.3 One-shot Training
@@
Next, we evaluate our approach in an extremely low data regime. Here, we train CoPL and CoCoOp with $1$ training instance per class for each of the image recognition task and test the accuracy on both seen and unseen classes within the dataset. In our experiments, we valuate across $7$ datasets, and discuss their performance on the seen and unseen classes next.
@@
Figure 4: Performance comparison of CoPL with CoCoOp for the seen class classification while trained on $1$-shot training instance per class.
@@
Figure 5: Performance comparison of CoPL with CoCoOp for the unseen classes classification while trained on $1$-shot training instance per class.
@@ 5 Discussions and Analysis
A key contribution of our research work is to utilize the localized image features for prompt learning. Further, we give weightage to the prompts based on the semantic relatedness. In this section, we critically analyse the contribution of using local features an opposed to global features. To evaluate this, we have conduct experiments on Caltech101 and DTD datasets. We selected these two dataset as they are primarily targeted for two different recognition tasks: object recognition and texture recognition. Similar to our earlier evaluation protocol, we evaluate the performance on the seen classes and unseen classes separately.
@@
Table 3: Ablation Study: Importance of aligning the local contextual image features with the prompt tokens
From Table 3, we can can clearly understand that for both seen and unseen classes, aligning local contextual features with prompts, helps CoPL to generalize better.
@@
Following the experimental analysis presented in CoCoOp [62], we evaluate the model efficacy where test dataset consists of both seen and unseen classes. In this case, during training the model weights are updated based on the seen classes but has not considered the unseen classes. Thus, during testing model will be performing zero-shot classification on unseen classes.
Methodology Accuracy
CLIP 65.2
CoOp 65.6
CoCoOp 69.1
CoPL 74.7
Table 4: Ablation study: average accuracy on $11$ datasets. The models are trained on training instances for the seen classes. Test datasets are the combination of both seen and unseen classes.
We observe in Table 4, that CoPL comfortably outperforms all the baseline models, suggesting that local image feature alignment with prompts help to generalize across all the classes within a dataset.
@@
We experimentally analyse and quantify the extra wall-clock time that is required for CoPL when compared to the closest baseline CoCoOp. We use Caltech101 for this experiment. For training $800$ data-points (corresponding to $50$ classes) , CoPL took $28$ minutes whereas evaluation on $1549$ data-points took $2$ minutes $05$ seconds. On the other hand CoCoOp takes $21$ minutes $35$ seconds for training and $1$ minute $34$ second for inference. From this, we clearly understand that CoPL significantly improves generalization without sacrificing too much on the computational overhead.
6 Limitation
@@
7 Conclusion
In this paper, we present CoPL: Contextual Prompt Learning which can align prompts to the corresponding contextual local image features. During alignment, we also produces a set of attention weights for the prompt vectors that are semantically related to local image regions. Extensive experimental evaluation on $11$ image recognition datasets showcases the efficacy of the methodology in understanding the semantic relationship between the images and the prompts. Moreover, the state-of-the-art zero-shot and few-shot results justify our claim of making CoPL better in generalization by aligning the local features of the images to prompts. The capability of outperforming CLIP in different settings justifies our argument of making dynamic prompt vectors comparable to human-annotated prompts, which opens up the floor to adapt the methodology for the various downstream tasks. In future, we plan to make COPL capable of understanding user intents to make local edits on the images.

### v2

@@ Introduction
We follow Zhou et al. (Zhou et al. 2022b) to evaluate our model on $11$ image classification dataset of varying complexity. The datasets include: generic classification datasets like ImageNet (Deng et al. 2009b) and Caltech-101 (Fei-Fei, Fergus, and Perona 2004); curated fine-grained datasets like OxfordPets (Parkhi et al. 2012), StanfordCars (Krause et al. 2013), Flowers102 (Nilsback and Zisserman 2008), Food101 (Bossard, Guillaumin, and Gool 2014) and FGVCAircraft (Maji et al. 2013); scene, action, texture and satellite image recognition datasets from SUN397 (Xiao et al. 2010), UCF101 (Soomro, Zamir, and Shah 2012), DTD (Cimpoi et al. 2014) and EuroSat (Helber et al. 2018) respectively. For the few-shot experiments, we follow Zhou et al. (Zhou et al. 2022a) to randomly sample datapoints for training. The models are evaluated on the entire test set to report the accuracy.
@@
In CLIP architecture, we use ViT-B/$16$ as image encoder. Our prompt token length is $4$. All our models are trained with a batch size $1$ for $10$ epochs on a single $16$ GB Tesla T4444 GPU system. Our starting learning rate is $0.002$ and used cosine learning rate scheduler. Our warm-up with a constant learning rate is $0.00001$.
@@
We compare our approach with CoCoOp (Zhou et al. 2022a), CoOp (Zhou et al. 2022b), KgCoOp (Yao, Zhang, and Xu 2023), ProGrad (Zhu et al. 2023) and also large-scale zero-shot methodology CLIP (Radford et al. 2021). While comparing with CLIP, we indeed compare our learned prompt embeddings with manually designed prompts. We closely follow most of our experimental setting with that of Zhou et al., (Zhou et al. 2022a).
@@
Table 2: Table shows the zero-shot image recognition results of CoPL across multiple dataset. We trained models on Caltech-101 dataset and tested on the other datasets. These results suggests very strong generalization ability of CoPL.
Models Source Target
ImageNet ImageNetV2 ImageNet-Sk ImageNet-A ImageNet-R Average
CLIP 66.7 60.8 46.1 47.7 73.9 57.1
CoCoOp 71.0 64.0 48.7 50.6 76.1 59.9
CoOp 71.5 64.2 47.9 49.7 75.2 59.2
ProGrad 72.2 64.7 47.6 49.3 74.5 59.0
KgCoOp 71.2 64.1 48.9 50.6 76.7 60.1
MaPLe 70.7 64.0 49.1 50.9 76.9 60.2
CoPrompt 70.8 64.2 49.4 50.5 77.5 60.4
CoPL 71.7 65.2 49.4 50.8 78.6 61.0
Table 3: Comparison of CoPL with existing approaches in domain generalization setting. On an average, CoPL outperforms other baseline results.
Knowledge Transfer to Unseen Classes
One of the main downsides of CoOp is that it is unable to generalize to the unseen classes while being trained on the base classes. The other methods improve over CoOp, but they miss out on capturing the local features of the image and do not give weightage to the semantically aligned learnable prompts, which is the key focus area of our work. Following the implementation of Zhou et al., (Zhou et al. 2022a), we conducted our experiments on the above mentioned $11$ datasets both for seen and unseen classes. While training is only conducted on the base classes, during testing we transfer the learnt knowledge to classify unseen classes as well as seen classes.
From Table 1, we observe that there is a decrease in performance for CoOp methodology over the zero-shot large-scale method CLIP for unseen classes. Though introduction of conditional prompting improve the performance of CoCoOp, in most of the cases it fails to generalize to unseen classes with the dynamically learnt prompts (in $7$ out of $11$ datasets CLIP outperform CoCoOp methodology). In fact KgCoOp, though it is not conditioned on the image features during prompt learning, outperforms CoCoOp in many cases.
Interestingly, our proposed approach CoPL, improves the accuracy of unseen classes over CoCoOp and CoOp for $10$ different tasks. It also outperforms KgCoOp and ProGrad during unseen class classification in $6$ tasks. Moreover, the harmonic mean for all $7$ tasks are greater than the current state-of-the-art model, which specifies that the learned prompts are more generalizable across domains and tasks. It is important to note that CoPL comfortably outperforms manual prompt based methodology CLIP on $5$ different tasks. With the relatively complex dataset like EuroSAT, where localization is hard to be aligned due to the nature of satellite imagery, for the seen classes CoPL outperformed CoCoOp by $1.8\%$, KgCoOp by $3.6\%$ and CLIP by $32.8\%$ on accuracy.
@@
In Figure 3, we further analyse the absolute performance gain on the unseen classes over CoOp and CoCoOp methodologies. When compared with CoOp on UFC101 dataset, CoPL improve the accuracy by $20.6\%$, showing the capability of the dynamic prompt vectors to learn semantic relatedness conditioned on the local image features. On FGVC-Aircraft dataset, the absolute performance gain of $9.0\%$ and $7.6\%$ for CoOp and CoCoOp respectively. Performance boost of $8.9\%$ on the DTD dataset, confirms the capability of COPL in identifying different textures, opening up the possibility to deploy the model for design understanding. CoPL outperforms CLIP by $2.7\%$ on the FGVC-Aircraft dataset, bringing out its ability to transfer the learnt knowledge to the relatively complex unseen classes.
@@
In general, for both seen and unseen dataset, CoPL outperforms CoCoOp by $2.7\%$, KgCoOp by $1.5\%$ and ProGrad by $2.4\%$ in accuracy. Also, it achieves a gain of $6.8\%$ in accuracy over manual prompt base methodology CLIP, suggesting that conditional alignment of prompts to local features of the image can generalize the model better to diverse recognition tasks. Ideally, the manually designed prompts should have been well aligned with the images, thus CLIP can be considered as the best performing model. It is interesting to note that CoPL, on an average across all $11$ datasets, achieves state-of-the-art performance by outperforming CLIP by $1.4\%$ in accuracy. This justifies our assertion of making dynamic prompt vectors more contextualized like human annotated prompts by observing the different local features of the image.
Inter Dataset Zero-shot Performance
@@
Here, we test the mettle of our approach on an even more challenging setting. We learn the model on a dataset, and evaluate the model on a different dataset to see how transferable the learned representations are. Concretely, we train the model on Caltech101 dataset and evaluate its performance on the rest of the datasets. Observe, in Table 2, CoPL outperforms CoCoOp across most of the datasets. For complex dataset like FGVCAircraft, CoPL gains $2.9\%$ Caltech101 being a general purpose object classification dataset, we observe impressive performance on OxfordPets and Food101 dataset. Interestingly, though trained on object classification dataset, CoPL is able to transfer the knowledge to do texture recognition task on DTD dataset, outperforming CoCoOp by $2.5\%$ on accuracy. In the most challenging EuroSAT dataset, consists of images taken from satellite, CoPL achieves $58.1\%$ accuracy, comparable to the performance of the CoCoOp. These highlights that CoPL able to learn the semantics of the images by identifying the localized features and align it with the learnable textual prompts.
Domain Generalization
Domain generalization matrix establishes the claim of a model being more adaptable to target dataset from same class but having different data distribution compared to sourced domain. Here with the existing baseline we have also compared our method with MaPLe (Khattak et al. 2023) and CoPrompt (Roy and Etemad 2023). We trained the model on Imagenet and tested the model on ImageNetV2, ImageNet-Sketch, ImageNet-A, and ImageNet-R. Observe, in Table 3, CoPL on an average outperforms all the baseline methodologies. This indicates that, by learning to focus on local image features, CoPL able to figure out the subjective image-prompt mapping, which makes CoPL more robust while working on out-of-distribution dataset.
One-shot Training
@@
We evaluate our approach in an extremely low data regime. Here, we train CoPL and CoCoOp with $1$ training instance per class for each of the image recognition task and test the accuracy on both seen and unseen classes within the dataset. In our experiments, we valuate across $7$ datasets, and discuss their performance on the seen and unseen classes next.
@@
Figure 4: Comparison of CoPL with CoCoOp for the seen class classification while trained on $1$-shot setting.
@@
Figure 5: Comparison of CoPL with CoCoOp for the unseen classes classification while trained on $1$-shot setting.
@@
A key contribution of CoPL is to utilize the localized image features for prompt learning. Further, we give weightage to the prompts based on the semantic relatedness. In this section, we critically analyse the contribution of using local features an opposed to global features. To evaluate this, we have conduct experiments on Caltech101 and DTD datasets. We selected these two dataset as they are primarily targeted for two different recognition tasks: object recognition and texture recognition. Similar to our earlier evaluation protocol, we evaluate the performance on the seen classes and unseen classes separately. From Table 4, we can can clearly understand that for both seen and unseen classes, aligning local contextual features with prompts, helps CoPL to generalize better.
@@
Table 4: Ablation Study: importance of aligning the local contextual image features with the prompt tokens.
@@
Following the experimental analysis presented in CoCoOp (Zhou et al. 2022a), we evaluate the model efficacy where test dataset consists of both seen and unseen classes. In this case, during training the model weights are updated based on the seen classes but has not considered the unseen classes. Thus, during testing model will be performing zero-shot classification on unseen classes.
Methodology CLIP CoOp CoCoOp CoPL
Accuracy 65.2 65.6 69.1 74.7
Table 5: Average accuracy on $11$ datasets, when the test set contains both the seen and unseen classes. The models are trained on training instances for the seen classes. CoPL optimally balances stability and plasticity.
We observe in Table 5, that CoPL comfortably outperforms all the baseline models, suggesting that local image feature alignment with prompts help to generalize better.
@@
We experimentally analyse and quantify the extra wall-clock time that is required for CoPL when compared to the closest baseline CoCoOp. We use Caltech101 for this experiment. For training $800$ data-points (corresponding to $50$ classes) , CoPL takes $28$ minutes whereas evaluation on $1549$ data-points took $2$ minutes $05$ seconds. On the other hand CoCoOp takes $21$ minutes $35$ seconds for training and $1$ minute $34$ second for inference. From this, we understand that CoPL improves generalization without sacrificing too much on the computational overhead.
Limitation
@@
Conclusion
In this paper, we present CoPL: Contextual Prompt Learning, which can align prompts to the corresponding contextual local image features. During alignment, we also produces a set of attention weights for the prompt vectors, that are semantically related to local image regions. Extensive experimental evaluation on $11$ image recognition datasets showcases the efficacy of CoPL in understanding the semantic relationship between the images and the prompts. Moreover, the state-of-the-art zero-shot and few-shot results justify our claim of making CoPL better in generalization by aligning the local features of the images to prompts. In future, we plan to make CoPL capable of understanding user intents to make local edits on images.